### v1

# PEVLM: Parallel Encoding for Vision-Language Models

Abstract
Vision-Language Models (VLMs) have demonstrated strong performance in video-language tasks, yet their application to long video understanding remains constrained by the quadratic complexity of standard attention mechanisms. In this paper, we propose PEVLM, a parallel encoding strategy specifically designed to improve the prefill efficiency of VLMs without requiring model finetuning. PEVLM partitions the input into block-wise segments with a shared sink, preserves Full-Attention positional embeddings, and aligns attention weights to mimic Full-Attention distributions. This design reduces attention computation from $O((T\times N)^{2})$ to $O(T\times N)$ while maintaining high accuracy. Extensive experiments on the LongVideoBench benchmark show that PEVLM achieves up to 8.37% accuracy improvement over existing inference-efficient methods and delivers up to 7.47x speedup in attention computation and 40% reduction in end-to-end latency. Under strict latency constraints, PEVLM significantly outperforms baselines, raising accuracy from 23.26% to 61.03%. These results highlight PEVLM’s effectiveness for low-latency, long-context video understanding, making it well-suited for real-world applications such as autonomous driving.
1 Background
1.1 Challenges in Long Context Processing
In the realm of computer vision and natural language processing integration, Vision-Language Models (VLMs) have demonstrated remarkable performance in various multimodal tasks, particularly in video-text understanding and generation (alayrac2022flamingo, ; radford2021learning, ; vlm_performance_example, ). However, their application to long video processing remains computationally challenging due to the quadratic complexity of self-attention mechanisms (long_video_processing_challenges_example, ; vaswani2017attention, ). Specifically, VLMs usually extract frame-wise features and integrate textual context information, leading to a drastic increase in context length as each frame is tokenized into multiple embeddings.
The high computational complexity of VLMs presents a significant barrier to their practical use and deployment in real-world scenarios (li2022video, ). Take autonomous driving as an example: training advanced autonomous driving models requires extensive long video data, which in turn demands efficient data pre-processing and annotation techniques. Traditional manual annotation methods and classical annotation models are inefficient in handling the huge volume of long video data. Although VLMs offer the possibility of automated annotation, the computational cost of processing long videos is prohibitively high. Moreover, autonomous driving systems impose strict requirements on low latency execution and high accuracy due to the critical need for real-time decision making to ensure safety (zhou2022efficient, ). However, on-board computing resources are severely limited, and the latency introduced by long-context attention mechanisms in VLMs significantly degrades the response time of autonomous systems. This situation highlights the urgent challenge of improving the inference performance of the VLMs with limited latency while maintaining its accuracy.
In autonomous driving scenarios, the system typically does not generate lengthy CoT (chain-of-thought) analyses, but instead needs to output only short environmental features and final decisions. This results in the prefill stage accounting for a substantial proportion of the overall inference time, becoming the primary performance bottleneck. However, most existing methods for optimizing the prefill stage involve a trade-off between performance improvement and accuracy degradation. Given that autonomous driving and other safety-critical applications are highly intolerant of accuracy loss, these methods are not applicable. Therefore, our research focuses on search for novel strategies to optimize the prefill stage latency of VLMs with minimal impact on accuracy.
1.2 Existing Solutions and Limitations
A widely adopted technique in LLMs is the parallel-encoding mechanism (li2024focusllmpreciseunderstandinglong, ; liu2024star, ; ma2025blockattentionefficientprefilling, ; ratner2023parallelcontextwindowslarge, ; yen2024longcontextlanguagemodelingparallel, ; yang2025ape, ). In this approach, the input context is divided into blocks, and each block is encoded into Key-Value (KV) states independently. This means that tokens from different blocks do not attend to each other during encodin. During decoding, however, the model computes attention using the complete KVCache that contains all historical blocks.
In addition, several block-based methods have been proposed that introduce training procedures to further enhance model accuracy and performance (ma2025blockattentionefficientprefilling, ; yuan2025nativesparseattentionhardwarealigned, ). However, not all application scenarios have sufficient resources to support large-scale training. On the other hand, many platforms offering inference services cannot control users’ preferences for model usage, nor can they mandate further fine-tuning on existing models. This is especially true for autonomous driving models that have undergone continuous training and iteration over months or even years. Therefore, we prioritize finetune-free approaches for optimization.
The parallel-encoding mechanism omits the attention calculation between blocks, reduces the computational complexity from $O(L^{2})$ to $O(L)$. It also mitigates the problem of "lost in the middle" (Liu2023LostIT, ; yang2025ape, ). Several architectural improvements have been proposed to enhance the parallel-encoding mechanism:
  • Adaptive Parallel Encoding (APE) improves prefill efficiency by introducing shared prefixes, an attention temperature parameter, and a scaling factor to better align the distribution of parallel-encoding with that of sequential encoding (yang2025ape, ). Although these adjustments lead to improved accuracy, they are highly sensitive to length, quantity, and content, making them difficult to tune and deploy in practice.
  • Star Attention, introduces a two-phase block-sparse approximation that improves computational efficiency by sharding attention across multiple hosts while minimizing communication overhead and reducing computational complexity (liu2024star, ). Although primarily designed for distributed settings, its core algorithms are mathematically similar to APE when temperature and scaling are not applied.
  • Block-Attention applies position re-encoding to allow reuse of prefilled contexts, and maintains model accuracy through fine-tuning (guu2020retrieval, ). Without fine-tuning, its accuracy suffers due to the absence of a Sink Block, though the idea of retaining Full-Attention position encodings proves insightful.
Thus, a natural research question thus arises: Can such methods be directly applied to Vision-Language Models (VLMs) to accelerate inference and deployment pipelines?
2 Our Observations and Contributions
In Table 1, we evaluate several existing methods and find that parallel-encoding-based approaches are also effective in VLMs. These methods achieve competitive accuracy on LongVideoBench (longvideobench2024, ) without requiring any modifications. This observation is consistent with findings from APE in LLM scenarios (yang2025ape, ).
2.1 Observations Specific to VLM
Despite the similarities between LLMs and VLMs, there are important distinctions. As shown in Figure 1: (i) In VLMs, attention sinks are present not only in the initial system prompt, but also in the early video frames, which aligns with the findings of Kang et al. (kang2025toldvisualattentionsink, ). To achieve optimal performance, it may be necessary to include all frames with high attention scores in the sink. This requirement may vary slightly across models; for the model shown in the figure (Qwen2.5-VL), the optimal range appears to be around 8–12 frames. (ii) Existing parallel-encoding-based are primarily designed for RAG (Retrieval-Augmented Generation) tasks. In such tasks, reusing the KVCache from historical passages is crucial for efficiency, and the order of passages often does not matter. Consequently, position embeddings are typically reused across blocks.
However, video-understanding tasks differ significantly: Maintaining the correct temporal order between video frames is essential. Reusing position embeddings indiscriminately across blocks can impair temporal comprehension of the model.
(a) Full-Attention
(b) Reuse Position
(c) Keep Full-Attention Position Embedding
Figure 1: Comparison of Attention Weight Distributions across Video Contexts. The top row shows the attention weights for all tokens, while the bottom row shows those for the video tokens only. (a) Full-Attention: The textual tokens and the initial video frames receive the highest attention weights, whereas the remaining video frames are assigned relatively lower attention weights. (b) Parallel Encoding with Reused Position Embeddings: The overall attention weights distribution resembles Full-Attention, with higher attention weights for the textual tokens and initial video frames. While compared to Full-Attention, the initial video frames receive even higher attention weights, and the remaining video frames exhibit more uniform and lower attention weights. (c) Parallel Encoding with Full-Attention Position Embedding: The attention pattern remains similar to (a) and (b), highlighting the textual tokens and the initial video frames. However, the attention weight distribution within the video frames more closely aligns with that of Full-Attention.
Therefore, as illustrated in Figure 1, we perform an attention weight analysis for VLMs and make two key findings: (i) Initial segments of both text and video receive high attention weights, emphasizing the need to include early video frames when selecting sink tokens. (ii) Using Full-Attention position embeddings (rather than reusing position embeddings across blocks) yields attention distributions more similar to Full-Attention, leading to better alignment with the model’s intended behavior. These insights guide the design of VLM-specific adaptations of parallel-encoding that preserve both efficiency and contextual integrity.
2.2 Our Contributions
Basing on these observations, we have conducted a series of studies and experiments. And proposed a novel attention mechanism for VLMs named Parallel Encoding for Vision-Language Models (PEVLM), which is a finetune-free method specifically designed to speed up prefill of VLM-based long-video processing. By alleviating computational bottlenecks in long video prefill, PEVLM facilitates the efficient deployment of VLMs in various practical application scenarios on both clouds and terminal equipments.
Our contributions involve:
  • We introduce PEVLM, a blockwise parallelization strategy for VLMs. PEVLM reduces the computational complexity of the attention mechanism from $O((T\times N)^{2})$ to $O(T\times N)$, which leading to a significant improvement in prefill efficiency.
  • We recover the accuracy of PEVLM with three alignment steps: (i) Segmenting contexts into blocks by video frames rather than tokens. (ii) Using the system promots and the initial frames of the video as Sink Block for all blocks to avoid the duplication of abnormal distribution of initial tokens. (iii) Preserving the position embeddings of Full-Attention instead of resuing position embeddings across blocks. This suggests that PEVLM serves as a performance optimization mechanism rather than endowing the model with the capability to process context lengths that exceed its inherent capacity.
  • We empirically show that (i) PEVLM gets better accuracy than Block-Attention, APE and Star Attention in long video processing tasks, achieving accuracy improvements ranging from 0.68% to 8.37% across different models. (ii) For 100k token contexts, the attention layer achieves a 7.47× speedup, while the end-to-end processing time is reduced by 40% without compromising generation quality. This highlights the practical viability and superiority of PEVLM in cloud deployment scenarios. (iii) Under fixed latency constraints, PEVLM increases accuracy from 23.26% to 61.03%, showcasing its critical value for latency sensitive applications (e.g., autonomous driving).
3 PEVLM
PEVLM is an advanced attention mechanism designed to improve computational efficiency of prefill while maintaining global contextual consistency. This approach draws from multiple existing techniques to address computational bottlenecks commonly encountered in long video processing.
3.1 Standard Attention Mechanism
In a standard Softmax attention, we attend the query to all past KV states using the following formula:
$O=Softmax(\frac{QK^{T}}{\sqrt{d}})V\ \ \ Q\in\mathbb{R}^{n\times d}\ \ \ K,V\in\mathbb{R}^{m\times d}$ (1)
where $Q$ is the query state, and $K$ and $V$ denote the key and value states, respectively.
Attention Sink. First discovered by StreamingLLM (mit2023streamingllm, ), that one or more tokens at the start position of an LLM receive higher attention scores. Some VLM-related studies also show that the attention sink phenomenon exists in VLMs but is more distributed among image tokens (kang2025toldvisualattentionsink, ).
Position Embedding. To effectively process sequential input, LLMs require position embeddings, such as absolute position embeddings (devlin-etal-2019-bert, )(vaswani2017attention, ) and relative position embeddings (press2022trainshorttestlong, )(SU2024127063, ). For VLMs, contextual positional information is more critical. In particular, for some VLM models like Qwen2.5-VL, their position embeddings also incorporate spatio-temporal information between video frames, making position embeddings increasingly essential for VLMs.
Computational Complexity. In standard Softmax attention mechanisms, the total OPs is
$OPs_{Attn}=2HL^{2},$ (2)
where $L$ denotes the context length and $H$ is the hidden size. So, the computational complexity is $O(L^{2})$. As $L$ increases, the computational cost increases significantly, restricting the inference speed and scalability. It is even worse for VLMs. Given a video with $T$ frames, where each frame is represented by $N$ tokens (typically, $N$ ranges between 200 and 600), the resulting context length $L$ scales to $(T\times N)$ and complexity to $O((T\times N)^{2})$, making standard attention mechanisms impractical for long video understanding tasks.
3.2 Parallel Encoding
3.2.1 Partitioning Strategy
Figure 2: The PEVLM Masks.
APE (yang2025ape, ) and Star-Attention (liu2024star, ) have demonstrated that sharing a common attention sink across blocks significantly enhances model accuracy. Building on recent advances in efficient attention mechanisms and long-context modeling (mit2023streamingllm, ; liu2024star, ; xiao2023efficient, ; yang2025ape, ), we propose a Sink Block mechanism to improve inference quality. As illustrated in Figure 2, our partitioning strategy consists of three core components:
  • Sink Block: The initial text tokens (e.g., BOS token and system prompts) and the first several video frames are grouped into a dedicated Sink Block.
  • Context Blocks: The remaining video content is uniformly partitioned into Context Blocks by frames to reduce computational overhead and enable parallelization.
  • Question Block: The text tokens that follow the video input are left unsegmented as Quesion Block.
Since the information within each video frame represents spatial content at a specific moment, while information across frames reflects temporal dynamics, videos naturally possess structural boundaries. If the video are divided into sink and context blocks by token count, this may compromise the spatial integrity of the boundary frames shared between adjacent blocks. So we divide the video by frame.
In addition, as analyzed in Section 2.1, the number of frames selected for the Sink Block depends primarily on the distribution of attention weights. We conduct further empirical studies and analyses on this in Section 5.2. On the other hand, the number of frames per Context Block involves a trade-off between accuracy and computational efficiency, which we investigate and analyze in Section 5.3.
3.2.2 Formulations
For different blocks, PEVLM applies distinct computation methods to ensure efficient processing and information integration:
$\text{Attn}_{s}\ =f(Q_{s},K_{s},V_{s}),$ (3)
$\text{Attn}_{c_{i}}=f(Q_{c_{i}},K_{s+c_{0}+\dots+c_{i-1}},V_{s+c_{0}+\dots+c_{i-1}}),$ (4)
$\text{Attn}_{q}\ =f(Q_{q},K_{s+c_{\text{all}}+q},V_{s+c_{\text{all}}+q}),$ (5)
where $s$ and $q$ are the Sink Block and question block, $c_{i}$ is the ith Context Blocks; $Attn_{s}$, $Attn_{c_{i}}$ and $Attn_{q}$ denote the attention of the $s$, $c_{i}$ and $q$. $f(\cdot)$ represents the standard Softmax attention in Equation (1), and $Q$, $K$ and $V$ denote the query, key, and value matrices.
The total OPs of PEVLM is
$OPs_{PEVLMAttn}=OPs_{SinkAttn}+N\times OPs_{BlockAttn}+OPs_{QuestAttn},$ (6)
$OPs_{SinkAttn}=2HS^{2},$ (7)
$OPs_{BlocksAttn}=2HB(S+B),$ (8)
$OPs_{QuestAttn}=2HQL,$ (9)
where $S$ denotes the sink size, $N$ denotes the block number, $B$ denotes the block size, $Q$ denotes the query size and $H$ is the hidden size. $L$ is the total token number and $L=S+N\times B+Q$.
After simplification:
$OPs_{PEVLMAttn}=2H(S^{2}+Q^{2}+NB^{2}+QS+NQB+NSB).$ (10)
As B is a static number, the computational complexity of PEVLM is $O(L)$.
3.2.3 Position Encoding in VLMs
Figure 3: Position Encoding in PEVLM
As described in Section 2.1, similar to the case in LLMs, applying parallel-encoding-based methods in VLMs suffers from the misalignment of attention weight distributions when compared to Full-Attention. However, introducing additional hyperparameters to adjust attention weights for better alignment is not ideal for efficient deployment. Therefore, inspired by Block Attention (ma2025blockattentionefficientprefilling, ), we retain Full-Attention Position Embedding as illustrated in Figure 3.
Due to maintaining the original position embedding, PEVLM cannot enable the model to handle longer video contexts beyond its inherent capabilities like some other parallel-encoding-based methods. For example, LongVILA-7B-256f can originally process 256 frames at most. Although using PEVLM can accelerate the inference speed, it still processes 256 frames at most. This is not a problem as our goal is to accelerate the prefill.
4 Experiments
In this section, we test PEVLM from multiple perspectives against several different VLMs. The primary objective is to evaluate the accuracy, computational efficiency, and real-world applicability of PEVLM for long-video processing.
4.1 Accuracy Evaluation
The main objective of this experiment is to analyze the impact of PEVLM on accuracy in long-video understanding tasks.
4.1.1 Setup
We have chosen the LongVideoBench (longvideobench2024, ) as a benchmark, which is designed to comprehensively evaluate large multimodal models on long-term multimodal understanding. We have carried out experiments using LLaVA-Video (zhang2024videoinstructiontuningsynthetic, ), Qwen2.5-VL (qwen2.5-VL, ), and LongVILA (chen2024longvilascalinglongcontextvisual, ) as test models. To minimize any adverse effects on model accuracy, all evaluations were performed using bf16 precision.
As shown in Table 1, we compared the accuracy of Full-Attention, Block-Attention, APE, Star-Attention, and PEVLM in the LongVideoBench dataset. Since the other methods are finetune-free, we did not finetune the weights in the Block Attention test. For APE and Star-Attention, we tested two scenarios: (i) reusing position encoding across all blocks; (ii) retaining the position encoding of Full-Attention. For PEVLM, we set the sink size and block size to about 16 frames, which corresponds to approximately 4k tokens for Qwen2.5-VL and LongVILA, and about 3k tokens for LLaVA-Video.
4.1.2 Results
Table 1: LongVideoBench Accuracy
model method sink(anchor)_size block_size Correct $\uparrow$
Full Attn - - 60.43%
Block Attn (wo finetune) 0 4096 54.23%
APE 256 4096 NaN
Qwen2.5-VL- APE (retain pos) 256 4096 60.21%
7B-Instruct Star Attn 4096 4096 NaN
Star Attn (retain pos) 4096 4096 61.41%
PEVLM sys+16frames 16frames 62.23%
Full Attn - - 52.21%
Block Attn (wo finetune) 0 4096 51.76%
APE 256 4096 51.68%
LongVILA- APE (retain pos) 256 4096 53.18%
7B-256f Star Attn 4096 4096 46.75%
Star Attn (retain pos) 4096 4096 53.63%
PEVLM sys+16frames 16frames 53.70%
Full Attn - - 59.99%
Block Attn (wo finetune) 0 4096 55.50%
APE 256 4096 58.41%
LLaVA-Video- APE (retain pos) 256 4096 58.49%
7B-Qwen2 (128) Star Attn 4096 4096 58.71%
Star Attn (retain pos) 4096 4096 58.49%
PEVLM sys+16frames 16frames 59.01%
The experiments demonstrate that PEVLM achieves an accuracy comparable to or even exceeding that of Full-Attention across the three tested models. Both APE and Star-Attention can also approach or surpass Full-Attention’s accuracy when retaining Full-Attention’s position encoding. However, for Block-Attention, due to the lack of resources for alignment fine-tuning, its accuracy shows a significant decline compared to Full-Attention.
Without keeping the position encoding of Full-Attention, the output of the Qwen2.5-VL-7B-Instruct model exhibited severe output garbling and semantic confusion, and the LongVILA model also experienced noticeable performance degradation. However, the LLaVA-Video model showed insensitivity to changes in position encoding, with negligible accuracy variation observed regardless of whether Full-Attention position encoding was retained.
4.2 Performance Evaluation
In this section, we evaluate the performance of PEVLM. To facilitate experimentation, we built PEVLM on the SGLang framework, which is a fast serving framework for large language models and vision-language models, widely used in cloud production deployment. And to better reflect practical deployment scenarios, we designed the experiments from two perspectives. (i) We first test with workloads of approximately 100k tokens to assess the acceleration performance of PEVLM in cloud computing environments. (ii) Furthermore, we evaluate PEVLM on the LongVideoBench dataset under a given latency constraint, in order to measure its effectiveness in edge scenarios where both computational resources and latency are limited.
4.2.1 Setup
Since the LLM component in VLM models remains largely consistent across different implementations, performance differences arise primarily from hyperparameter configurations rather than model weights. Consequently, we conducted computational efficiency evaluations on the Qwen2.5-VL model. The primary objective of this experiment is to assess the improvements in computational efficiency introduced by PEVLM in computational efficiency. All experiments were carried out on an H20-96G machine. Given that PEVLM primarily optimizes the attention mechanism within the LLM, we measured the execution time of both the entire LLM module and its individual attention layers separately.
Figure 4: We evaluated the performance of PEVLM under different configurations of (sink_size / block_size). Using the Qwen2.5-VL-7B-Instruct model, we tested its LLM and Attention time on a video dataset with an average duration of 10 minutes (approximately 106k tokens per sample). "Attn Time" designates the Attention Computation time, whereas "Other Time" signifies the aggregate of all remaining computational costs within the LLM excluding Attention Computation. The "Attn SpeedUP" represents the performance acceleration of PEVLM relative to the "Full Attn". The "LLM SpeedUP" denotes the overall performance acceleration of the LLM in PEVLM compared to the "Full Attn".
4.2.2 Results
As illustrated in Figure 4, the end-to-end inference speed of the LLM is improved by 2.58× to 3.14×. Notably, the attention module achieves a speedup ranging from 7.79× to 33.81×, while the execution time of non-attention components remains largely unaffected. As the block size or sink size increases, the acceleration effect of the attention module decreases significantly. This observation is consistent with the analysis presented in the Appendix. Moreover, because non-attention components constitute a substantial portion of the total runtime in PEVLM, the overall end-to-end speedup of the LLM moderately decreases from 3.14× to 2.58×.
4.3 Accuracy with Limited Latency
One key challenge in deploying VLMs on edge devices is the limited computational resources and strict latency requirements. To evaluate the acceleration performance of PEVLM under this scenario, we designed a simulation accordingly. A simplified experimental setup was implemented to assess the efficiency of PEVLM in edge deployment settings.
4.3.1 Setup
We use LongVideoBench as the evaluation benchmark and introduce a latency constraint during inference: any sample whose inference time exceeds the predefined latency threshold is considered a failure. The experiments are conducted on the H20-96B platform. We select Qwen2.5-VL-7B-Instruct with Full-Attention as the baseline, and compare several methods, including Full-Attention, Block Attention, APE, Star Attention, PEVLM, and using smaller models (Qwen2.5-VL-3B-Instruct). Since the other methods are finetune-free, we did not finetune the weights in the Block Attention test. Although position encoding does not affect inference lantency, reusing position encoding performs extremely poorly accuracy on the Qwen2.5-VL model. Therefore, in the APE and Star Attention tests, we adopted the method of retaining Full-Attention’s position encoding.
4.3.2 Results
Table 2: LongVideoBench Accuracy with Limited Latency
method sink(anchor) block No Limit $\uparrow$ 40s $\uparrow$ 30s $\uparrow$ 20s $\uparrow$
Full Attn - - 60.43% 59.69% 27.08% 23.26%
Full Attn (3B) - - 55.05% 55.05% 55.05% 24.38%
Block Attn 0 4096 54.00% 54.00% 53.78% 53.55%
APE 256 4096 59.91% 59.91% 59.91% 59.61%
Star Attn 4096 4096 61.41% 61.41% 61.41% 28.20%
PEVLM sys+2frames 16frames 61.18% 61.18% 61.18% 61.03%
PEVLM sys+16frames 16frames 62.23% 62.23% 62.23% 60.28%
As shown in Table 2, the accuracy of Full-Attention drops significantly when the latency constraint is reduced from 40s to 30s, while the accuracy of Full-Attention 3B also declines sharply when the latency constraint is tightened from 30s to 20s. Due to its larger sink and block sizes, StarAttn gets more OPs and exhibits a noticeable accuracy drop between 30s and 20s. In contrast, Block Attention, APE, and PEVLM maintain relatively high accuracy even at 20s. Notably, although larger sink and block sizes (e.g., sys+16frames, block=16frames) achieve higher accuracy under looser latency constraints, the smaller sink size (sink=sys+2frames, block=16frames) delivers better accuracy at the 20s latency limit due to its lower latency.
5 How does each component in PEVLM contribute to the performance?
5.1 Split The Video by Frames instead of Tokens
Due to the inherent nature of videos being composed of multiple frames, where intra-frame information captures spatial details at a single time point, and inter-frame information represents temporal dynamics, ignoring this structure and arbitrarily dividing blocks based on token count may compromise the spatial integrity at critical temporal boundaries. Therefore, we choose to divide the video into blocks based on frames. This design is supported by the comparison between Star Attention with Full-Attention positional encoding and PEVLM, as shown in Table 1. Across different models, frame-based block partitioning consistently provides accuracy gains, albeit varying in magnitude, over token-based partitioning.
5.2 Sink Size
As analyzed in Section 3.2.2, the inference latency of PEVLM increases quadratically with the sink size, which is consistent with the experimental results shown in Figure 4.
Furthermore, we investigate the impact of sink size on model accuracy. As observed in Section 2.1, it is essential to involve the initial frames of the video in the sink. This observation is validated in Figure 5, where incorporating the initial frames into the sink leads to varying degrees of accuracy improvement across different models.
Moreover, as shown in Figure 5, the accuracy of Qwen2.5-VL and LLaVA-Video shows an upward trend as the sink size increases, though the improvement remains marginal. In contrast, LongVILA exhibits fluctuations in accuracy with varying sink sizes but does not show a consistent upward or downward trend. However, because of the quadratic increase in latency, using longer sink sizes is not favorable for performance. Therefore, it may be necessary to conduct a simple test prior to deployment to determine an appropriate sink size. For the three models shown in Figure 5, setting the sink size between 14 and 18 already yields strong accuracy. If latency constraints are also taken into account, this could lead to a trade-off between performance and accuracy. For example, as demonstrated in the results of Section 4.3, under a 20-second latency constraint, using only sys+2frames as the Sink Block achieves better accuracy than using sys+16frames.
Figure 5: The impact of sink/block sizes on accuracy
5.3 Block Size
As analyzed in Section 3.2.2, the inference latency of PEVLM also increases quadratically with the block size as well as the sink size, which is consistent with the experimental results shown in Figure 4. However, the block size has a more significant impact on performance because the coefficient of block-size’s quadratic term in Equation 10 is much larger than that of the sink size.
We investigate the impact of block size on model accuracy here. As shown in Figure 5, the model accuracy exhibits an upward trend as the block size increases, similar to sink size. However, due to the quadratic growth in latency, a trade-off arises between computational efficiency and accuracy.
5.4 Position Encoding
As analyzed in Section 2.1, using Full-Attention position embeddings, rather than reusing position embeddings across blocks, should yield attention distributions that more closely resemble those of Full-Attention, thereby aligning better with the model’s intended behavior. This is validated by the experiments presented in Section 4.1. As shown in Table 1, for several parallel-encoding-based methods (Block-Attention, APE, Star Attention, and PEVLM), using Full-Attention position embeddings significantly improves accuracy in both Qwen2.5-VL and LongVILA. In contrast, for the LLaVA-Video model, the benefit is marginal, likely because its attention weight distribution remains relatively stable even when using parallel encoding.
6 Conclusion
This study investigates the potential of parallel-encoding-based methods in VLMs, which can reduce the computational complexity of attention mechanisms for fast inference and reuse positions for long-context processing but often result in performance degradation. To address this issue, we propose PEVLM, a training-free method designed to enable accurate, fast, and long-context video processing. PEVLM achieves this by aligning the attention weight distribution of parallel encoding with Full-Attention through two key steps: selecting an appropriate sink size and block size, and utilizing Full-Attention position embeddings. Empirically, we demonstrate that PEVLM enhances both accuracy and efficiency in long video processing tasks while successfully scaling to process hundreds of chunks in parallel across various settings.
7 Future Work
Although PEVLM has demonstrated substantial improvements in computational efficiency and reasoning performance for long video processing tasks, there remains ample opportunity for further research and enhancement. In future work, we plan to explore the following directions.
In autonomous driving scenarios, a common strategy is to use several historical frames as input to make decisions for the next action. Due to the computational nature of the attention mechanism, most current approaches require re-computation of all historical frames as the temporal step advances. Building upon the current PEVLM framework, we propose integrating the rotating positional embedding method from Block-Attention (ma2025blockattentionefficientprefilling, ), which enables the reuse of the KVCache from all historical frames preceding the current timestep. This approach has the potential to yield an order-of-magnitude improvement in efficiency for such applications.
Many of the existing methods based on parallel-encoding, including PEVLM, can achieve accuracy that exceeds Full-Attention. This phenomenon is attributed to the fact that parallel-encoding-based methods can mitigate the "lost in the middle" problem, thereby enabling them to achieve accuracy surpassing Full-Attention when processing long contexts (yang2025ape, ). The reason why parallel-encoding can mitigate the "lost in the middle" issue is hypothesized to be due to its shortening of the distance between the context and the query. However, more concrete evidence is needed to support this claim.